The length of a rectangle is three times its width. If the area of the rectangle is 192 cm^2 , find its perimeter.

Answer by nerdybill(5404)

The length of a rectangle is three times its width. If the area of the rectangle is 192 cm^2 , find its perimeter.
.
Let w = width
then
3w = length
.
192 = w(3w)
192 = 3w^2
64 = w^2
8 cm = w
.
length:
3w = 3*8 = 24 cm
.
Therefore perimeter is:
2(8+24) = 2(32) = 64 cm